# Improving threshold for fault-tolerant color code quantum computing by flagged weight optimization

### v1

@@ III Improving threshold for fault-tolerant color code quantum computing
Using conventional syndrome measurement circuits and weights results in a lower threshold in color codes due to their high-weight stabilizer generators. Here, we propose a method to improve the thresholds of color codes under the circuit-level noise. In the following, we first describe the syndrome measurement circuit employed here. Subsequently, we explain how to set the weights of decoders. The deflagging procedure, which is carried out to further improve the performance, is also introduced. We also propose a method for estimating conditional error probabilities, which is also one of our main contributions. Following that, the results of the numerical simulation we performed are presented.
@@ III.1 Syndrome measurement gadget
In a syndrome measurement, we use a flag gadget for each face as an ancilla instead of using a single ancilla qubit for each face. Fig. 8 shows two types of flag gadgets used for $X$ stabilizer measurements: a two-qubit flag gadget [22] and a four-qubit flag gadget [40]. In these gadgets, the qubits prepared in $\ket{+}$ act as syndrome qubits, while those prepared in |$\ket{0}$ act as flag qubits. Flag gadgets for $Z$ stabilizer measurements are obtained by reversing the basis and the direction of the CNOT gates.
@@
After preparing the cat state, they interact with data qubits and finally perform post-processing and measurements. The cat state allows us to parallelize CNOT gates and reduce the depth for the syndrome measurement, resulting in less idling noise. Additionally, a decrease in the error propagation paths leads to a reduction in the error propagation. For the (4.8.8) color code, the four-qubit flag gadget is used for the weight-8 stabilizer measurements, and the two-qubit flag gadget is used for the weight-four stabilizer measurements. We show the CNOT order for the (4.8.8) color code when using the flag gadgets in Fig. 9. The CNOT gates are applied to the two neighboring data qubits from each qubit in the gadget. For the (6.6.6) color code, the two-qubit flag gadget is used for each stabilizer measurement. The CNOT order for the (6.6.6) color code is shown in Fig. 10. The CNOT gates are applied to the three neighboring data qubits from each qubit in the gadget for the weight-6 stabilizer measurements and to the two neighboring data qubits for the weight-four stabilizer measurements. Since an ancilla qubit interacts with only two or three data qubits, the propagation of errors from the ancilla qubits to the data qubits is suppressed compared to the conventional syndrome measurement circuit. Furthermore, the depth of the CNOT gates acting on the data qubits has been reduced to three steps in both lattices, which is the minimum in the case of color codes. This is because in color codes, up to three stabilizer generators are involved for a data qubit, and at each time, at most one operation can be applied to any qubit. The $X$ stabilizer measurement circuits for each face of the (4.8.8) color code and the (6.6.6) color code are shown in Figs. 11 and 12, respectively.
Figure 9: CNOT order for the (4.8.8) color code with the flag gadgets. The total number of qubits required for this method is $(5d^{2}+4d-5)/4$.
Figure 10: CNOT order for the (6.6.6) color code with the flag gadgets. The total number of qubits required for this method is $(3d^{2}-1)/2$.
@@ III.2 Flagged weight optimization
We explain the setting of the decoder weights. Here, each cycle of a syndrome measurement consists of performing an $X$ stabilizer measurement followed by a $Z$ stabilizer measurement, as shown in Fig. 13. Let us consider the case of performing $X$ error correction. If an $X$ error occurred in ancilla qubits for an $X$ stabilizer measurement propagates to data qubits as a hook error, it is detected in the subsequent $Z$ stabilizer measurement. Thus, the flag values of the $X$ stabilizer measurement at time $t$ provide information about the $X$ data errors at time $t$. Here, in order to incorporate the flag information into the decoder weights defined by Eqs. (12) and (13), we set the weights using conditional error probabilities conditioned on the flag values. For weights of $X$ data errors, we use the conditional error probabilities conditioned on all flag values in the measurements of the $X$ stabilizers acting on the corresponding data qubits at the same time step:
$w^{(t)}_{v}=-\log\frac{p\left(x^{(t)}_{v}=1|\bigcup_{f:v\in f}\mathcal{F}^{(t)}_{f,X}\right)}{1-p\left(x^{(t)}_{v}=1|\bigcup_{f:v\in f}\mathcal{F}^{(t)}_{f,X}\right)}.$ (15)
Here, $\mathcal{F}^{(t)}_{f,\sigma}$ is a set of flag values in the flag gadget used for measuring the $\sigma\in\{X,Z\}$ stabilizers defined on the face $f$ at time $t$. Flag values of a $Z$ stabilizer measurement explicitly provide information only about $Z$ errors occured in the ancilla qubits, but they also implicitly provide information about $X$ errors occured in the ancilla qubits. Therefore, for weights of measurement errors, we use the conditional error probabilities conditioned on the flag values in the measurements of the $Z$ stabilizers that measure the corresponding syndrome:
@@
On the other hand, in the case of performing $Z$ error correction, the weights for $Z$ data errors and measurement errors are set as follows:
$w^{(t)}_{v}=\begin{cases}-\log\frac{p\left(z^{(1)}_{v}=1\right)}{1-p\left(z^{(1)}_{v}=1\right)}&\text{if }t=1,\\ -\log\frac{p\left(z^{(t)}_{v}=1|\bigcup_{f:v\in f}\mathcal{F}^{(t-1)}_{f,Z}\right)}{1-p\left(z^{(t)}_{v}=1|\bigcup_{f:v\in f}\mathcal{F}^{(t-1)}_{f,Z}\right)}&\text{otherwise.},\end{cases}$ (17)
@@
In Eq. (17), we do not use flag information for the weight of the $Z$ data errors at the initial time step. This is because the procedure for a syndrome measurement consists of a cycle of first performing an $X$ stabilizer measurement, followed by a $Z$ stabilizer measurement, so there is no flag information for the $Z$ data errors at the initial time step.
@@ III.3 Deflagging
In addition to the flagged weight optimization, we perform a deflagging procedure [32, 33], i.e., immediately apply proper Pauli operators to the data qubits based on the flag values. Without directly applying the Pauli operators to the data qubits, we can perform the same operation by just recording the Pauli operators as classical information and updating the Pauli frame [41, 42]. Deflagging is also performed when estimating the conditional error probabilities. This deflagging procedure enables us to correct some errors that cannot be corrected with only the flagged weight optimization. For the details of the deflagging procedure and how it improves logical error rates, see Appendix B.
@@ III.4 Estimating conditional error probabilities
The conditional error probabilities in Eqs. (15)-(18) need to be estimated before decoding. Here, we propose a method to estimate the conditional error probabilities that is accurate, efficient, and can be used even when the underlying noise is a priori unknown. We estimate the conditional error probabilities by running a tailored quantum circuit multiple times. We employ quantum circuits $C_{X}$ and $C_{Z}$, each tailored for estimating the conditional error probabilities used in the weights of the $X$ and $Z$ error correction decoders, respectively. These quantum circuits are modified versions of a one-cycle syndrome measurement circuit, where the initial states of certain qubits are altered and the data qubits are measured transversally at the end. In $C_{X}$, an $X$ stabilizer measurement circuit is executed with all data qubits and ancilla qubits initialized to |$\ket{0}$, followed by an original $Z$ stabilizer measurement. Then, the data qubits are measured transversally in the $Z$-basis. In the case of $C_{Z}$, we use a one-cycle syndrome measurement circuit, where a $Z$ stabilizer measurement is performed, followed by an $X$ stabilizer measurement. The initial states of all data qubits and ancilla qubits used for the $Z$ stabilizer measurement are prepared as $\ket{+}$. At the end, the data qubits are measured in the $X$-basis in a transversal manner. These circuits enable the detection of data errors, and also allow us to estimate measurement errors occurred in the circuits. We show $C_{X}$ for the case of the distance-three (4.8.8) color code in Fig. 14.
Figure 14: The quantum circuit $C_{X}$ for the distance-three (4.8.8) color code. This circuit is tailored for estimating the conditional error probabilities used in the weights of the $X$ error correction decoder.
@@
    Repeat the steps up to this point a sufficient number of times, and finally divide the total count of data errors and measurement errors by the number of times each flag value pattern is measured.
Since the syndrome measurement is translationally invariant, the conditional error probabilities estimated by this procedure is used at all time steps. Note that in $C_{X}$ and $C_{Z}$, preparing the initial state in this way is necessary to detect data errors. This is because the measurement outcomes become random when the data qubits are directly measured in the original syndrome measurement circuit. This randomness arises from two factors: first, the data qubits are in a superposition state; and second, a certain stabilizer operator is always applied to the data qubits due to the entangled ancilla qubits. Once the conditional error probabilities are estimated offline, those values can be used in all subsequent decoding.
@@ IV.1 Settings
We perform Monte Carlo simulations to compute the logical error rates when the proposed method is applied to the integer programming decoder under the circuit-level noise. The number of samples in the Monte Carlo simulations is $10^{6}$. Since the quantum state after the decoding and recovery operation in space-time may not necessarily be in the code space, we perform an ideal error correction at the final time-slice [20]. For comparison, we also compute the logical error rates of the existing method that uses a single ancilla qubit for each syndrome measurement and employs uniform decoder weights. We call this existing method the single ancilla method in the following. We use Stim [43] to implement the quantum circuits and CPLEX [44] for the integer programming solver.
@@
$p_{\mathrm{L}}=c\left(\frac{p}{p_{\mathrm{th}}}\right)^{\alpha\left(\frac{d+1}{2}\right)}.$ (19)
Here, $c$ is a constant, $p$ is a physical error rate, $p_{\mathrm{th}}$ is a threshold, $d$ is a code distance, and $\alpha$ is a constant indicating the effective code distance. We fit the logical error rates using Eq. (19) with $c$, $p_{\mathrm{th}}$, and $\alpha$ as fitting parameters.
@@ IV.2 Results
Figure 15: Logical error rates of the (4.8.8) color code for various code distances. (a) Logical $X$ error rates. (b) Logical $Z$ error rates.
Figure 16: Logical error rates of the (6.6.6) color code for various code distances. (a) Logical $X$ error rates. (b) Logical $Z$ error rates.
@@
Logical $X$ and $Z$ error rates for each physical error rate are shown in Fig. 15 for the (4.8.8) color code, and in Fig. 16 for the (6.6.6) color code. For the fitting, we use the data in the region $p\in[10^{-4},10^{-3}]$ for $d=7$ and $d=9$. The values obtained by fitting the logical $X$ and $Z$ error rates of the (4.8.8) color code in Fig. 15 are shown in Tables 1 and 2, respectively. For the (6.6.6) color code, the values obtained by fitting the logical $X$ and $Z$ error rates in Fig. 16 are shown in Tables 3 and 4, respectively. As shown in Table 1, the $X$ error threshold of the proposed method for the (4.8.8) color code is 0.29(1)%, which is almost 1.9 times higher than that of the single ancilla method. From Table 2, it can be seen that the $Z$ error threshold of the proposed method for the (4.8.8) color code is 0.268(9)%, which is almost 1.8 times higher than that of the single ancilla method. The $Z$ error threshold is slightly lower than the $X$ error threshold, mainly because there is no flag information for the $Z$ data errors in the initial time step. Thus, the $Z$ error threshold of 0.268(9)sets the overall threshold. As shown in Tables 3 and 4, the $X$ and $Z$ error thresholds of the proposed method for the (6.6.6) color code are 0.37(1)and 0.363(9)%, respectively. The $X$ error threshold is almost 1.4 times higher, and the $Z$ error threshold is almost 1.3 times higher than that of the single ancilla method, respectively. Here, note that the threshold obtained from the fitting in the region of the low logical error rates and the threshold derived from the intersection of the logical error rates at different code distances are slightly different. In the case of the (4.8.8) color code, except for the $X$ error threshold of the proposed method, the latter value is higher. For the (6.6.6) color code, the former value is higher in any case. In terms of the effective code distance, the proposed method improves $\alpha$ in either case. Thereby, in both Figs. 15 and 16, the logical error rates are almost one order of magnitude lower than those of the single ancilla method when the physical error rate is low, i.e., around $p=10^{-4}$.
@@
Compared to the thresholds reported in the previous studies, for the (4.8.8) color code, we surpassed the thresholds of all previous studies [20, 21]. For the (6.6.6) color code, we achieved the threshold almost the same as the highest threshold of 0.37(1)[23], among the previous studies that adopt the same noise model. Note that in Ref. [24], a threshold of around 0.47% was achieved, but Ref. [24] adopts a noise model that assumes fewer errors in state preparation and measurement than the noise model we adopt. It should also be noted that a high threshold does not necessarily guarantee lower logical error rates far below the threshold, because the logical error rates are also influenced by the effective code distance.
@@
$c$ $p_{\mathrm{th}}$ $\alpha$ Proposed method 0.13(2) 0.0029(1) 0.69(1) Single ancilla method 0.049(4) 0.00154(6) 0.537(8)
@@
$c$ $p_{\mathrm{th}}$ $\alpha$ Proposed method 0.11(1) 0.00268(9) 0.679(9) Single ancilla method 0.046(3) 0.00148(5) 0.535(6)
@@
$c$ $p_{\mathrm{th}}$ $\alpha$ Proposed method 0.12(1) 0.0037(1) 0.72(1) Single ancilla method 0.116(6) 0.00274(6) 0.610(5)
@@
$c$ $p_{\mathrm{th}}$ $\alpha$ Proposed method 0.117(7) 0.00363(9) 0.675(6) Single ancilla method 0.111(6) 0.00270(6) 0.600(5)

### v2

@@ III Improving threshold for fault-tolerant color code quantum computing
Using conventional syndrome measurement circuits and weights results in a lower threshold in color codes due to their high-weight stabilizer generators. Here, we propose a method to improve the thresholds of color codes under the circuit-level noise. In the following, we first describe the syndrome measurement circuit employed here. Subsequently, we explain how to set the weights of decoders. The deflagging procedure, which is carried out to further improve the performance, is also introduced. We also propose a method for estimating conditional error probabilities, which is also one of our main contributions.
@@ III.1 Syndrome measurement gadget
In a syndrome measurement, we use a flag gadget for each face to extract a syndrome instead of using a single ancilla qubit for each face. Fig. 8 shows two types of flag gadgets used for $X$ stabilizer measurements: a two-qubit flag gadget [25] and a four-qubit flag gadget [47]. In these gadgets, the qubits prepared in $\ket{+}$ act as syndrome qubits, while those prepared in |$\ket{0}$ act as flag qubits. Flag gadgets for $Z$ stabilizer measurements are obtained by reversing the basis and the direction of the CNOT gates.
@@
After preparing the cat state, they interact with data qubits and finally perform post-processing and measurements. The cat state allows us to parallelize CNOT gates and reduce the depth for the syndrome measurement, resulting in less idling noise. Additionally, a decrease in the error propagation paths leads to a reduction in the error propagation. For the (4.8.8) color code, the four-qubit flag gadget is used for the weight-8 stabilizer measurements, and the two-qubit flag gadget is used for the weight-four stabilizer measurements, the same as in Ref. [24]. We show the CNOT order for the (4.8.8) color code when using the flag gadgets in Fig. 9. The CNOT gates are applied to the two neighboring data qubits from each qubit in the gadget. For the (6.6.6) color code, the two-qubit flag gadget is used for each stabilizer measurement. The CNOT order for the (6.6.6) color code is shown in Fig. 10. The CNOT gates are applied to the three neighboring data qubits from each qubit in the gadget for the weight-6 stabilizer measurements and to the two neighboring data qubits for the weight-four stabilizer measurements. Since an ancilla qubit interacts with only two or three data qubits, the propagation of errors from the ancilla qubits to the data qubits is suppressed compared to the conventional syndrome measurement circuit. Furthermore, the depth of the CNOT gates acting on the data qubits has been reduced to three steps in both lattices, which is the minimum in the case of color codes. This is because in color codes, up to three stabilizer generators are involved for a data qubit, and at each time, at most one operation can be applied to any qubit. The total number of data and ancilla qubits required for this way of measuring syndrome is $n_{4.8.8}^{\text{FWO}}(d)=(5d^{2}+4d-5)/4$ for the (4.8.8) color code and $n_{6.6.6}^{\text{FWO}}(d)=(3d^{2}-1)/2$ for the (6.6.6) color code. The $X$ stabilizer measurement circuits for each face of the (4.8.8) color code and the (6.6.6) color code are shown in Figs. 11 and 12, respectively.
Figure 9: CNOT order for the (4.8.8) color code with the flag gadgets. The total number of qubits required for this method is $n_{4.8.8}^{\text{FWO}}(d)=(5d^{2}+4d-5)/4$.
Figure 10: CNOT order for the (6.6.6) color code with the flag gadgets. The total number of qubits required for this method is $n_{6.6.6}^{\text{FWO}}(d)=(3d^{2}-1)/2$.
@@ III.2 Flagged weight optimization
We explain the setting of the decoder weights. Here, each cycle of a syndrome measurement consists of performing an $X$ stabilizer measurement followed by a $Z$ stabilizer measurement, as shown in Fig. 13. Let us consider the case of performing $X$ error correction. If an $X$ error occurring on an ancilla qubit for an $X$ stabilizer measurement propagates to data qubits as a hook error, it is detected in the subsequent $Z$ stabilizer measurement. Thus, the flag values of the $X$ stabilizer measurement at time $t$ provide information about the $X$ data errors at time $t$. Note that flag values are not perfect, as errors can also occur on the flag qubits. Here, in order to incorporate the flag information into the decoder weights defined by Eqs. (9) and (10), we set the weights using conditional error probabilities conditioned on the flag values. For weights of edges corresponding to $X$ data errors, we use the conditional error probabilities conditioned on all flag values in the measurements of the $X$ stabilizers acting on the corresponding data qubits at the same time step:
$w^{(t)}_{v}=-\log\frac{p\left(x^{(t)}_{v}=1|\bigcup_{f:v\in\delta f}\mathcal{F}^{(t)}_{f,X}\right)}{1-p\left(x^{(t)}_{v}=1|\bigcup_{f:v\in\delta f}\mathcal{F}^{(t)}_{f,X}\right)}.$ (12)
Here, $\mathcal{F}^{(t)}_{f,\sigma}$ is a set of flag values in the flag gadget used for measuring the $\sigma\in\{X,Z\}$ stabilizers defined on the face $f$ at time $t$. Flag values of a $Z$ stabilizer measurement explicitly provide information only about $Z$ errors occurred in the ancilla qubits, but they also implicitly provide information about $X$ errors occurred in the ancilla qubits. To elaborate, a trivial (i.e., unflipped) flag value implies that either an $X$ error that is not correlated with a $Z$ error occurred or no error occurred at a certain location. Therefore, for weights of edges corresponding to measurement errors, we use the conditional error probabilities conditioned on the flag values in the measurements of the $Z$ stabilizers that measure the corresponding syndrome:
@@
On the other hand, in the case of performing $Z$ error correction, the weights of edges corresponding to $Z$ data errors and measurement errors are set as follows:
$w^{(t)}_{v}=\begin{cases}-\log\frac{p\left(z^{(1)}_{v}=1\right)}{1-p\left(z^{(1)}_{v}=1\right)}&\text{if }t=1,\\ -\log\frac{p\left(z^{(t)}_{v}=1|\bigcup_{f:v\in\delta f}\mathcal{F}^{(t-1)}_{f,Z}\right)}{1-p\left(z^{(t)}_{v}=1|\bigcup_{f:v\in\delta f}\mathcal{F}^{(t-1)}_{f,Z}\right)}&\text{otherwise.},\end{cases}$ (14)
@@
In Eq. (14), we do not use flag information for the weights of the edges corresponding to $Z$ data errors at the initial time step. This is because the procedure for a syndrome measurement consists of a cycle of first performing an $X$ stabilizer measurement, followed by a $Z$ stabilizer measurement, so there is no flag information for the $Z$ data errors at the initial time step.
@@ III.3 Deflagging
In addition to FWO, we perform a deflagging procedure proposed in Refs. [38, 39], i.e., immediately apply Pauli operators corresponding to the errors that are implied by the flag values to the data qubits. Without directly applying the Pauli operators to the data qubits, we can perform the same operation by just recording the Pauli operators as classical information and updating the Pauli frame [48, 49]. Deflagging is also performed when estimating the conditional error probabilities. This deflagging procedure enables us to correct some errors that cannot be corrected with only FWO. For the details of the deflagging procedure and how it improves logical error rates, see Appendix B.
@@ III.4 Estimating conditional error probabilities
The conditional error probabilities in Eqs. (12)-(15) need to be estimated before decoding. Here, we propose a method to estimate the conditional error probabilities that is accurate, efficient, and can be used even when the underlying noise is a priori unknown. We estimate the conditional error probabilities by running a tailored quantum circuit multiple times. We employ quantum circuits $C_{X}$ and $C_{Z}$, each tailored for estimating the conditional error probabilities used in the edge weights in the $X$ and $Z$ error correction decoders, respectively. These quantum circuits are modified versions of a one-cycle syndrome measurement circuit, where the initial states of certain qubits are altered and the data qubits are measured transversally at the end. In $C_{X}$, an $X$ stabilizer measurement circuit is executed with all data qubits and ancilla qubits initialized to |$\ket{0}$, followed by a $Z$ stabilizer measurement. Then, the data qubits are measured transversally in the $Z$-basis. In the case of $C_{Z}$, we use a one-cycle syndrome measurement circuit, where a $Z$ stabilizer measurement is performed, followed by an $X$ stabilizer measurement. The initial states of all data qubits and ancilla qubits used for the $Z$ stabilizer measurement are prepared as $\ket{+}$. At the end, the data qubits are measured in the $X$-basis in a transversal manner. These circuits enable a direct detection of data errors and also allow us to estimate measurement errors that can occur in the circuits. Here, a direct detection means detecting errors occurring in a qubit solely based on the outcome of single-qubit measurement for that qubit. We show $C_{X}$ for the case of the distance-three (4.8.8) color code in Fig. 14.
Figure 14: The quantum circuit $C_{X}$ for the distance-three (4.8.8) color code. This circuit is tailored for estimating the conditional error probabilities used in the weights in the $X$ error correction decoder.
@@
    Repeat the steps up to this point a sufficiently large number of times. Then, to estimate the probability of a data or measurement error given an observed set of flag outcomes, divide the total number of times each error was observed with the set of flag outcomes, by the total number of times the set of flag outcomes was observed.
Since the syndrome measurement is repetitive, the conditional error probabilities estimated by this procedure is used at all time steps. Note that in $C_{X}$ and $C_{Z}$, preparing the initial state in this way is necessary for the direct detection of data errors. This is because without such a modification of the initial states, each measurement outcome of the transversal measurement for data qubits at the end of the syndrome measurement circuit cannot be used for the direct detection of data errors. The reason is that, when detecting $X$ ($Z$) errors, the data qubits and the ancilla qubits used in the $X$ ($Z$) stabilizer measurement circuits are in a superposition state. Once the conditional error probabilities are estimated offline, those values can be used in all subsequent decoding.
@@ IV.1 Settings
We perform Monte Carlo simulations to compute the logical error rates when the proposed method is applied to the integer programming decoder under the circuit-level noise. The number of samples in the Monte Carlo simulations is $10^{6}$. In order to evaluate logical error rates, we need to compare the input logical state with the output logical state. Since the quantum state after the decoding and recovery operation in space-time may not necessarily be in the code space, we perform an ideal error correction at the final time-slice [49, 23] to project the state onto the code space. For comparison, we also compute the logical error rates of the existing method that uses a single ancilla qubit for each syndrome measurement and employs uniform decoder weights. We call this existing method the single ancilla method in the following. We use Stim [50] to implement the quantum circuits and CPLEX [51] for the integer programming solver. The source code we used is available on GitHub [52].
@@
$p_{\mathrm{L}}=c\left(\frac{p}{p_{\mathrm{th}}^{*}}\right)^{\alpha\left(\frac{d+1}{2}\right)}.$ (16)
Here, $c$ is a constant, $p$ is a physical error rate, $p_{\mathrm{th}}^{*}$ is a threshold, $d$ is a code distance, and $\alpha$ is a constant indicating the effective code distance. We refer to the threshold $p_{\mathrm{th}}^{*}$ obtained from Eq. (16) as the scaling threshold. We fit the logical error rates using Eq. (16) with $c$, $p_{\mathrm{th}}^{*}$, and $\alpha$ as fitting parameters.
We also calculate the threshold obtained from the crossing point of data for different code distances. For a sufficiently large code distance $d$, the logical error rates $p_{L}$ around the threshold under circuit-level noise are expected to behave as follows [23, 53]:
$p_{\mathrm{L}}=A+B(p-p_{\mathrm{th}}^{\times})d^{1/\nu_{0}}+C(p-p_{\mathrm{th}}^{\times})^{2}d^{2/\nu_{0}}.$ (17)
Here, $A$, $B$, and $C$ are constants, $p$ is a physical error rate, $p_{\mathrm{th}}^{\times}$ is a threshold, $d$ is a code distance, and $\nu_{0}$ is a critical exponent. We refer to the threshold $p_{\mathrm{th}}^{\times}$ obtained from Eq. (17) as the cross threshold. We fit the logical error rates using Eq. (17) with $A$, $B$, $C$, $p_{\mathrm{th}}^{\times}$, and $\nu_{0}$ as fitting parameters.
@@ IV.2 Results
Figure 15: Logical error rates of the (4.8.8) color code for various code distances. The syndrome measurements are performed for $d$ rounds. The cross thresholds indicated by the gray vertical dashed lines in the figures are calculated by using the data of $d=7$ and $d=9$. (a) Logical $X$ error rates. The scaling threshold and cross threshold for the proposed method (solid lines) are $p_{\mathrm{th}}^{*}=0.29(1)\%$ and $p_{\mathrm{th}}^{\times}=0.279(1)\%$, respectively. The thresholds for the single ancilla method (dotted lines) are $p_{\mathrm{th}}^{*}=0.16(1)\%$ and $p_{\mathrm{th}}^{\times}=0.1694(5)\%$. (b) Logical $Z$ error rates. The thresholds for the proposed method (solid lines) are $p_{\mathrm{th}}^{*}=0.268(9)\%$ and $p_{\mathrm{th}}^{\times}=0.276(1)\%$. The thresholds for the single ancilla method (dotted lines) are $p_{\mathrm{th}}^{*}=0.14(1)\%$ and $p_{\mathrm{th}}^{\times}=0.164(2)\%$.
Figure 16: Logical error rates of the (6.6.6) color code for various code distances. The syndrome measurements are repeated $d$ times. The gray vertical dashed lines indicate the cross thresholds, which are calculated by using the data of $d=7$ and $d=9$. (a) Logical $X$ error rates. The scaling threshold and cross threshold for the proposed method (solid lines) are $p_{\mathrm{th}}^{*}=0.37(1)\%$ and $p_{\mathrm{th}}^{\times}=0.3575(6)\%$, respectively. The thresholds for the single ancilla method (dotted lines) are $p_{\mathrm{th}}^{*}=0.274(6)\%$ and $p_{\mathrm{th}}^{\times}=0.2547(8)\%$. (b) Logical $Z$ error rates. The thresholds for the proposed method (solid lines) are $p_{\mathrm{th}}^{*}=0.363(9)\%$ and $p_{\mathrm{th}}^{\times}=0.352(1)\%$. The thresholds for the single ancilla method (dotted lines) are $p_{\mathrm{th}}^{*}=0.270(6)\%$ and $p_{\mathrm{th}}^{\times}=0.2528(8)\%$.
@@
Logical $X$ and $Z$ error rates for each physical error rate are shown in Fig. 15 for the (4.8.8) color code, and in Fig. 16 for the (6.6.6) color code. For each, we fit the logical error rates with Eqs. (16) and (17) using the data for $d=7$ and $d=9$. When we fit them using Eq. (16), we use the data with a physical error rate $p$ well below the threshold, that is, in the range of $p\lessapprox p_{\mathrm{th}}^{\times}/2$ [27]. We use the data in the range of $p\in[10^{-4},10^{-3}]$ except for the single ancilla method for the (4.8.8) color code, where the range is $p\in[10^{-4},5\times 10^{-4}]$ so that the data is considered to be well below the threshold. When we fit the logical error rates using Eq. (17), we used five data points around the threshold. The values obtained by fitting the logical $X$ and $Z$ error rates of the (4.8.8) color code in Fig. 15 are shown in Tables 1 and 2, respectively. For the (6.6.6) color code, the values obtained by fitting the logical $X$ and $Z$ error rates in Fig. 16 are shown in Tables 3 and 4, respectively. As shown in Table 1, the $X$ error scaling threshold of the proposed method for the (4.8.8) color code is $p_{\mathrm{th}}^{*}$=0.29(1)%, which is almost 1.8 times higher than that of the single ancilla method. From Table 2, it can be seen that the $Z$ error scaling threshold of the proposed method for the (4.8.8) color code is $p_{\mathrm{th}}^{*}$=0.268(9)%, which is almost 1.9 times higher than that of the single ancilla method. The $Z$ error scaling threshold is slightly lower than the $X$ error scaling threshold, mainly because there is no flag information for the $Z$ data errors in the initial time step. Thus, the $Z$ error scaling threshold of $p_{\mathrm{th}}^{*}$=0.268(9)sets the overall scaling threshold. As shown in Tables 3 and 4, the $X$ and $Z$ error scaling thresholds of the proposed method for the (6.6.6) color code are $p_{\mathrm{th}}^{*}$=0.37(1)and $p_{\mathrm{th}}^{*}$=0.363(9)%, respectively. The $X$ error scaling threshold is almost 1.4 times higher, and the $Z$ error scaling threshold is almost 1.3 times higher than that of the single ancilla method, respectively. In all cases, the obtained cross threshold is nearly the same as the scaling threshold.
@@
Compared to the thresholds reported in the previous studies, for the (4.8.8) color code, we surpassed the thresholds of all previous studies [23, 24]. For the (6.6.6) color code, we achieved the threshold that is the same within statistical errors as the highest threshold of 0.37(1)[26], among the previous studies that adopt the same noise model. Note that in Ref. [28], a cross threshold of around 0.47% was achieved, but Ref. [28] adopts a noise model that assumes fewer errors in state preparation and measurement than the noise model we adopt. Ref. [27] obtained the cross threshold of around 0.46%, but the obtained scaling threshold is around 0.32%, which is lower than our value of 0.36%. It should also be noted that a high threshold does not necessarily guarantee lower logical error rates far below the threshold, because the logical error rates are also influenced by the effective code distance.
@@
$c$ $\alpha$ $p_{\mathrm{th}}^{*}$ $p_{\mathrm{th}}^{\times}$ Proposed 0.13(2) 0.69(1) 0.0029(1) 0.00279(1) Single ancilla 0.035(6) 0.47(1) 0.0016(1) 0.001694(5)
@@
$c$ $\alpha$ $p_{\mathrm{th}}^{*}$ $p_{\mathrm{th}}^{\times}$ Proposed 0.11(1) 0.679(9) 0.00268(9) 0.00276(1) Single ancilla 0.031(5) 0.49(1) 0.0014(1) 0.00164(2)
In terms of the effective code distance, the proposed method improves $\alpha$ in either case. Thereby, in both Figs. 15 and 16, the logical error rates are almost one order of magnitude lower than those of the single ancilla method for the same code distances when the physical error rate is low, i.e., around $p=10^{-4}$. Nonetheless, when the number of available physical qubits is given, comparing logical error rates of each method for the same code distance is not always a fair comparison. For the range of code distances we used for the simulation, i.e., $d\in\{5,7,9\}$, the total number of data and ancilla qubits satisfies the following relationships:
$n_{4.8.8}^{\text{FWO}}(d-2)<n_{4.8.8}^{\text{single}}(d)<n_{4.8.8}^{\text{FWO}}(d),$ (18)
$n_{6.6.6}^{\text{FWO}}(d-2)<n_{6.6.6}^{\text{single}}(d)<n_{6.6.6}^{\text{FWO}}(d).$ (19)
Comparing logical error rates of each method with the same code distance $d$ is justified only when the number of available physical qubits is assumed to be $n^{\prime}$ such that $n_{4.8.8}^{\text{FWO}}(d)<n^{\prime}<n_{4.8.8}^{\text{single}}(d+2)$ ($n_{6.6.6}^{\text{FWO}}(d)<n^{\prime}<n_{6.6.6}^{\text{single}}(d+2)$). However, in situations where $n^{\prime}$ physical qubits such that $n_{4.8.8}^{\text{single}}(d)<n^{\prime}<n_{4.8.8}^{\text{FWO}}(d)$ ($n_{6.6.6}^{\text{single}}(d)<n^{\prime}<n_{6.6.6}^{\text{FWO}}(d)$) are allowed to be used, the logical error rates of the single ancilla method with distance $d$ should be compared with the proposed method with distance $d-2$. From Fig. 15, it can be seen that the logical error rates of the proposed method with distance 5 (7) are lower than those of the single ancilla method with distance 7 (9) for the (4.8.8) color code, which means our proposed method improves the logical error rates compared to the single ancilla method even if the number of available qubits is limited to any specific number. In the case of the (6.6.6) color code, when the physical error rate is $p=10^{-4}$, the logical error rate of the proposed method with distance 5 is slightly higher than that of the single ancilla method with distance 7. However, in other physical error rates regimes, the logical error rates of the proposed method with distance 5 (7) are lower than or same as those of the single ancilla method with distance 7 (9).
@@
$c$ $\alpha$ $p_{\mathrm{th}}^{*}$ $p_{\mathrm{th}}^{\times}$ Proposed 0.12(1) 0.72(1) 0.0037(1) 0.003575(6) Single ancilla 0.116(6) 0.610(5) 0.00274(6) 0.002547(8)
@@
$c$ $\alpha$ $p_{\mathrm{th}}^{*}$ $p_{\mathrm{th}}^{\times}$ Proposed 0.117(7) 0.675(6) 0.00363(9) 0.00352(1) Single ancilla 0.111(6) 0.600(5) 0.00270(6) 0.002528(8)